ezarius03:

The system has the solution (3, –1), where F, G, H, P, Q, and R are nonzero real numbers. Select all the systems that are also guaranteed to have the solution (3, –1).

dude:

Uh you might want to include the "systems" it is talking about

ezarius03:

The system Px + Qy=R Fx + Gy=H has the solution (3, –1), where F, G, H, P, Q, and R are nonzero real numbers. Select all the systems that are also guaranteed to have the solution (3, –1). A (P + F)x + (Q + G)y = R + H Fx + Gy = H B (P + F)x + Qy = R + H Fx + (G + Q)y = H C Px + Qy = R (3P + F)x + (3Q + G)y = 3H+R D Px + Qy = R (F – 2P)x + (G – 2Q)y = H – 2R E Px + Qy = R 5Fx + 5Gy = 5H 14581 Px Qy R Fx Gy H

BigMac:

Very hard to read. Hard to tell which equations belong to which system if you only separate everything with one line.

sillybilly123: